# Learning to Steer Learners in Games

### v1

@@ 1 Introduction
To address this problem, the dominant approach to learning in games has been formulating it as an adversarial online learning problem. In this framing, to handle opponents with unknown objectives and learning rules, each player assumes they are faced with an arbitrary (and potentially adversarial) sequence of payoffs and seeks to find an algorithm that maximizes their own utility. Given this setup, a natural class of algorithms to choose from is the class of no-regret algorithms, which guarantees asymptotically optimal performance compared to the best fixed action in hindsight Cesa-Bianchi and Lugosi (2006). If all players use no-regret algorithms, it is well known that the average action of the players over the entire time horizon converges to a (coarse) correlated equilibrium Foster and Vohra (1998); Hart and Mas-Colell (2000). However, since the opponents are also learners rather than adversaries, adopting a no-regret algorithm may not be optimal.
@@
This problem has been well studied in recent years—though primarily in the case where the payoff matrix, and thus the objective of the opponent player, is known. Under such assumptions, it was shown in Deng et al. (2019) that if one player (the optimizer) deviates from using a no-regret algorithm, it can (under mild assumptions about the game instance) guarantee an asymptotic average payoff that is arbitrarily close to the Stackelberg value of the game by steering the no-regret player (the learner) to the Stackelberg equilibrium of the underlying matrix game. This is the highest attainable value if the optimizer plays one fixed mixed strategy. Further studies have focused on analyzing the steerability of smaller classes of algorithms, such as no-swap-regret Brown et al. (2023) or mean-based algorithms Arunachaleswaran et al. (2024), providing more insights into steering no-regret learners. Crucially, all these works assume knowledge of the learner’s objectives—knowledge that the optimizer itself may not even have.
@@
What information does the optimizer need in order to learn to achieve this goal?
@@ 2 Related Works
The problem of steering a no-regret learner in a repeated game has been the focus of several recent works, though often under strong assumptions on what information is available to the optimizer. Assuming known learner payoffs, Braverman et al. (2017) first introduced the problem of steering a learner in an auction setting. Subsequently, Deng et al. (2019) showed that the optimizer can guarantee at least the Stackelberg value against the learner in bimatrix games and Assos et al. (2024) studied the problem of utility maximization under the additional assumptions on the learner’s algorithm. Brown et al. (2023) also focused on smaller classes of no-regret algorithms such as no-swap-regret, anytime no-regret and no-adaptive-regret algorithms. While sharing a similar goal with our work, all of these works use the known learner payoff to design steering algorithms, which is not available in our setting. Without the knowledge of learner payoff, Brânzei et al. (2024) showed the steerability of the learner in a cake-cutting model, which can be viewed as a 2D special case of our problem. There is also a broader line of works regarding more general properties of interacting with no-regret learners, including Zhang et al. (2024) that considered the steering problem through direct payments to the learner and Arunachaleswaran et al. (2024) that studied the problem of pareto-optimality in the space of learning algorithms.
@@
The problem of steering is closely related to the problem of learning to play a Stackelberg equilibrium through repeated interactions. Thus, a particularly related line of work involves learning in unknown repeated Stackelberg games, where the decisions are made sequentially in each round. This problem has been well-studied in its own right but often under simplifying assumptions on the games or the responses of the follower (the learner in our framing of the problem). Letchford et al. (2009) and Peng et al. (2019) proposed algorithms for learning through interaction with a myopic best-responding agent, and Haghtalab et al. (2022) extended this framework to non-myopic agents with discounted utilities over time— a different setup from the one we consider. Other works have analyzed similar problems under different assumptions on the underlying game. In the control literature, Lauffer et al. (2022) studied the problem of learning in dynamic Stackelberg games and showed that one could learn the Stackelberg equilibrium. Maheshwari et al. (2024) studied a similar problem of learning Stackelberg equilibria in the context of continuous games; Goktas et al. (2022) studied the behavior of no-regret learning in a smaller class of zero-sum Stackelberg games, and the problem of steering learning agents has also emerged in the literature on strategic classification Zrnic et al. (2021). In each of these problems, the additional structure introduced into the games simplifies the task of learning the Stackelberg equilibrium through e.g., convexity or smoothness of the underlying optimization problem. Unfortunately, in bimatrix games, the Stackelberg optimization problem is both highly non-convex and discontinuous, vastly complicating the task of learning Stackelberg equilibria. Learning in Stackelberg games has also been studied in various application areas including security Blum et al. (2014); Balcan et al. (2015), calibration Haghtalab et al. (2023) and learning with side information Harris et al. (2024). Most results proposed in these works assume the follower uses (nearly) myopic best response dynamics, which does not fit into our setting.
@@
Several other relevant works do not directly fit into either ‘steering’ or ‘learning’. Conitzer and Sandholm (2006) first proposed an efficient algorithm of computing Stackelberg equilibria when the game is known through solving and combining several small linear programs, motivating our steering approach based on payoff matrix recovery. Gan et al. (2023) studied the notion of robust Stackelberg equilibrium allowing the follower to respond with any $\delta$-optimal response. Despite being formulated differently, this shares the same high-level idea with our approach of inducing pessimism to steer learners. Collina et al. (2024) considered the setting of finitely repeated Stackelberg games where the leader commitments and follower actions are adaptive to the gameplay history and proposed an efficient algorithm for approximating Stackelberg equilibria in the space of adaptive game playing algorithms.
@@ 3.2 Regret and No-regret Algorithms
Given some function $f:\mathbb{N}\rightarrow\mathbb{R}$ such that $f(T)=o(T)$ and an optimizer action sequence $\{x_{t}\}_{t=1}^{T}$, an interaction sequence $\{x_{t},y_{t}\}_{t=1}^{T}$ is $f$-no-regret for the learner if
@@
for some constant $C$, a learner algorithm $\mathcal{A}_{2}$ is $f$-no-regret on $\{x_{t}\}_{t=1}^{T}$ if
@@ 4 Impossibility of Learning to Steer Agents Who Use General No-regret Algorithms
Theorem 1 suggests that even if the optimizer knows that the learner payoff is one of the two different candidates $B_{1}$ or $B_{2}$, whatever algorithm $\mathcal{A}_{1}$ they try to come up with, there exists a no-regret algorithm $\mathcal{A}_{2}$ for the learner that can induce an $\Theta(T)$ Stackelberg regret to the optimizer in one of the two game instances. This suggests that we cannot hope to design a no-Stackelberg-regret algorithm for the optimizer that works simultaneously well on all game instances without any additional assumption on the learner besides it being no-regret. Interestingly, Theorem 6 in Brown et al. (2023) suggests that the optimizer is able to learn and steer a no-adaptive-regret learner, indicating the fundamental difference of learning to steer learners with different algorithm classes.
@@ 5.1 Facets and Equivalence Classes of Payoff Matrices
while similar definitions are made in Letchford et al. (2009); Peng et al. (2019), here we use the word facet because each $E_{i}$ induced by any matrix $B$ is indeed a polytope that is a subset of $\Delta_{m}$ and $\cup_{i\in[n]}E_{i}=\Delta_{m}$. Intuitively, the boundary of a facet specifies the critical hyperplane in the space of mixed strategies $\Delta_{m}$ of the optimizer, where the learner is indifferent between two (or more) pure strategies. In general, a Stackelberg equilibrium strategy $x^{*}$ stays at an extreme point of one facet, and therefore, in order to extract Stackelberg from the learner, the optimizer must be able to (or potentially implicitly) reconstruct the facet boundaries around the equilibrium point.
@@
At a Stackelberg equilibrium the learner must be indifferent between multiple pure strategies, which means that switching from the equilibrium pure strategy $y^{*}$ to another indifferent pure strategy $y^{\prime}$ does not incur additional regret to the learner, while potentially vastly degrading the optimizer’s payoff. Therefore, instead of simply selecting the equilibrium point, the optimizer must choose a pessimistic equilibrium point that sacrifices some utility from Stackelberg value to guarantee the learner responds with $y^{*}$. We define pessimistic facets as:
@@
As shown in Proposition 1, as long as the optimizer knows a complete set of $d$-pessimistic facets for each $i\in[n]$, it is able to compute an approximate Stackelberg equilibrium up to an error at the scale of $d$. We now extend this result to its ability to extract Stackelberg value against no-regret learners:
@@
To describe an equivalence class $\mathcal{B}$, observe that for all matrices $B$ in $\mathcal{B}$, the matrix $\mathcal{B}_{i}^{\circ}\in\mathbb{R}^{m\times(n-1)}$ defined by $(\mathcal{B}_{i}^{\circ})_{:,k}=(B_{:,k}-B_{:,i})/\max_{j_{1},j_{2}}\|B_{:,j_{1}}-B_{:,j_{2}}\|_{\infty}$ in each column will be the same for any fixed index $i\in[m]$ with the convention that $/0=0$ 0, so we can use $\mathcal{B}_{i}^{\circ}$ to represent the entire equivalence class. Based on this, we can also define the difference between two equivalence classes:
@@ 5.2 Steering with Payoff Class Estimation
We can see from Definition 11 that $C_{i}\leq 0$ if and only if (15) is feasible. We show in Section C.2 that if $C_{i}>0$ there exists a gap where (15) can be relaxed while still being feasible. If $C_{i}=0$, however, Definition 11 indicates that $\forall x,\max_{j}x^{T}\mathcal{B}_{i}^{\circ}(e_{j}-e_{i})\geq 0$, and $\exists x\in\Delta_{m},j\in[n]$ such that $x^{T}B_{i}(e_{j}-e_{i})=0$. Under this case, the facet $E_{i}(\mathcal{B})$ has zero volume and as long as the estimation $\hat{\mathcal{B}}$ is not precise, the facet $E_{i}^{-}(\hat{\mathcal{B}},d)$ could always be empty. Also, even if the optimizer knows the real underlying $\mathcal{B}$, since $e_{i}$ is weakly dominated, the learner is not steerable if $e_{i}$ happens to be the Stackelberg equilibrium since the learner is indifferent between $e_{i}$ and $e_{j}$. To avoid this special case (which occurs with probability 0 for uniformly randomly generated $B$ matrices Von Stengel and Zamir (2010)), we make the following assumption, as is standard (see e.g., Gan et al. (2023); Deng et al. (2019); Brown et al. (2023)):
@@ 6.1 Learning to Steer Ascending Learners with $n=2$
Suppose $m=n=2$ and the payoff matrix $B$ does not contain identical columns. If either one facet is empty, or each facet has diameter at least $d$ and $P_{2}$ uses an ascent algorithm $\mathcal{A}_{2}$ that is $f$-no-regret, Algorithm 1 with accuracy margin $d$ achieves a Stackelberg regret of at most $O(\frac{f(T)}{d}+dT-\log d)$ as long as $d=\Omega(\exp(-f(T)))$.

### v2

@@ 1 Introduction
To address this problem, the dominant approach to learning in games has been formulating it as an adversarial online learning problem. In this framing, to handle opponents with unknown objectives and learning rules, each player assumes they are faced with an arbitrary (and potentially adversarial) sequence of payoffs and seeks to find an algorithm that maximizes their own utility. Given this setup, a natural class of algorithms to choose from is the class of no-regret algorithms, which guarantees asymptotically optimal performance compared to the best fixed action in hindsight (Cesa-Bianchi and Lugosi, 2006). If all players use no-regret algorithms, it is well known that the average action of the players over the entire time horizon converges to a (coarse) correlated equilibrium (Foster and Vohra, 1998; Hart and Mas-Colell, 2000). However, since the opponents are also learners rather than adversaries, adopting a no-regret algorithm may not be optimal.
@@
This problem has been well studied in recent years—though primarily in the case where the payoff matrix, and thus the objective of the opponent player, is known. Under such assumptions, it was shown in (Deng et al., 2019) that if one player (the optimizer) deviates from using a no-regret algorithm, it can (under mild assumptions about the game instance) guarantee an asymptotic average payoff that is arbitrarily close to the Stackelberg value of the game by steering the no-regret player (the learner) to the Stackelberg equilibrium of the underlying matrix game. This is the highest attainable value if the optimizer plays one fixed mixed strategy. Further studies have focused on analyzing the steerability of smaller classes of algorithms, such as no-swap-regret (Brown et al., 2023) or mean-based algorithms (Arunachaleswaran et al., 2024), providing more insights into steering no-regret learners. Crucially, all these works assume knowledge of the learner’s objectives—knowledge that the optimizer itself may not even have.
@@
What information does the optimizer need in order to achieve this goal?
@@ 2 Related Works
The problem of steering a no-regret learner in a repeated game has been the focus of several recent works, though often under strong assumptions on what information is available to the optimizer. Assuming known learner payoffs, Braverman et al. (2017) first introduced the problem of steering a learner in an auction setting. Subsequently, Deng et al. (2019) showed that the optimizer can guarantee at least the Stackelberg value against the learner in bimatrix games and Assos et al. (2024) studied the problem of utility maximization under the additional assumptions on the learner’s algorithm. Brown et al. (2023) also focused on smaller classes of no-regret algorithms such as no-swap-regret, anytime no-regret and no-adaptive-regret algorithms. While sharing a similar goal with our work, all of these works use the known learner payoff to design steering algorithms, which is not available in our setting. Without the knowledge of learner payoff, Brânzei et al. (2024) showed the steerability of the learner in a cake-cutting model, which can be viewed as a 2D special case of our problem. Lin and Chen (2024) proposed a principal-agent framework and studied the optimal average utility that can be obtained by the optimizer assuming either a no-regret or a no-swap-regret algorithm under known learner payoff. There is also a broader line of works regarding more general properties of interacting with no-regret learners, including (Zhang et al., 2024) that considered the steering problem through direct payments to the learner and (Arunachaleswaran et al., 2024) that studied the problem of pareto-optimality in the space of learning algorithms.
@@
The problem of steering is closely related to the problem of learning to play a Stackelberg equilibrium through repeated interactions. Thus, a particularly related line of work involves learning in unknown repeated Stackelberg games, where the decisions are made sequentially in each round. This problem has been well-studied in its own right but often under simplifying assumptions on the games or the responses of the follower (the learner in our framing of the problem). Letchford et al. (2009) and Peng et al. (2019) proposed algorithms for learning through interaction with a myopic best-responding agent, and Haghtalab et al. (2022) extended this framework to non-myopic agents with discounted utilities over time— a different setup from the one we consider. Other works have analyzed similar problems under different assumptions on the underlying game. In the control literature, Lauffer et al. (2022) studied the problem of learning in dynamic Stackelberg games and showed that one could learn the Stackelberg equilibrium. Maheshwari et al. (2024) studied a similar problem of learning Stackelberg equilibria in the context of continuous games; Goktas et al. (2022) studied the behavior of no-regret learning in a smaller class of zero-sum Stackelberg games, and the problem of steering learning agents has also emerged in the literature on strategic classification (Zrnic et al., 2021). In each of these problems, the additional structure introduced into the games simplifies the task of learning the Stackelberg equilibrium through e.g., convexity or smoothness of the underlying optimization problem. Unfortunately, in bimatrix games, the Stackelberg optimization problem is both highly non-convex and discontinuous, vastly complicating the task of learning Stackelberg equilibria. Learning in Stackelberg games has also been studied in various application areas including security (Blum et al., 2014; Balcan et al., 2015), calibration (Haghtalab et al., 2023) and learning with side information (Harris et al., 2024). Most results proposed in these works assume the follower uses (nearly) myopic best response dynamics, which does not fit into our setting.
@@
Several other relevant works do not directly fit into either ‘steering’ or ‘learning’. Conitzer and Sandholm (2006) first proposed an efficient algorithm of computing Stackelberg equilibria when the game is known through solving and combining several small linear programs, motivating our steering approach based on payoff matrix recovery. Gan et al. (2023) studied the notion of robust Stackelberg equilibrium allowing the follower to respond with any $\delta$-optimal response. Despite being formulated differently, this shares the same high-level idea with our approach of inducing pessimism to steer learners. Collina et al. (2024) considered the setting of finitely repeated Stackelberg games where the leader commitments and follower actions are adaptive to the gameplay history and proposed an efficient algorithm for approximating Stackelberg equilibria in the space of adaptive game playing algorithms. There are also works from an empirical perspective that considers (and leverages) the learning behavior of other agents to achieve higher payoff and more stable learning process (Foerster et al., 2017; Lu et al., 2022).
@@ 3.2 Regret and No-regret Algorithms
Given some function $f:\mathbb{N}\rightarrow\mathbb{R}$ such that $f(T)=o(T)$ and an optimizer action sequence $\{x_{t}\}_{t=1}^{T}$, an interaction sequence $\{x_{t},y_{t}\}_{t=1}^{T}$ is $f$-no-regret for the learner (with constant $C$) if
@@
for some constant $C$, a learner algorithm $\mathcal{A}_{2}$ is $f$-no-regret (with constant $C$) on $\{x_{t}\}_{t=1}^{T}$ if
@@ 4 Impossibility of Learning to Steer Agents Who Use General No-regret Algorithms
The proof of Theorem 1 is deferred to Appendix A.
Theorem 1 suggests that even if the optimizer knows that the learner payoff is one of the two different candidates $B_{1}$ or $B_{2}$, whatever algorithm $\mathcal{A}_{1}$ they try to come up with, there exists a no-regret algorithm $\mathcal{A}_{2}$ for the learner that can induce an $\Theta(T)$ Stackelberg regret to the optimizer in one of the two game instances. The proof of Theorem 1 relies on first designing a game instance such that $G_{1}$ and $G_{2}$ has different Stackelberg equilibrium, and use a simple algorithm $\mathcal{A}_{2}^{\prime}$ against which the optimizer is not able to distinguish whether the realized learner payoff matrix is $B_{1}$ or $B_{2}$, before finally modifying it to be no-regret. The idea of constructing a pair of game instances that have different equilibrium but the same payoff function for one player is also used in the proof of Theorem 3 in (Bajaj et al., 2024), in which they showed that in a repeated two-player game where the opponent strategy is not known, no algorithm can achieve a bounded competitive ratio against itself (used by the opponent) for all such game instances.
This suggests that we cannot hope to design a no-Stackelberg-regret algorithm for the optimizer that works simultaneously well on all game instances without any additional assumption on the learner besides it being no-regret. Interestingly, Theorem 6 in (Brown et al., 2023) suggests that the optimizer is able to learn and steer a no-adaptive-regret learner, indicating the fundamental difference of learning to steer learners with different algorithm classes.
@@ 5.1 Facets and Equivalence Classes of Payoff Matrices
Intuitively, the boundary of a facet specifies the critical hyperplane in the space of mixed strategies $\Delta_{m}$ of the optimizer, where the learner is indifferent between two (or more) pure strategies. In general, a Stackelberg equilibrium strategy $x^{*}$ stays at an extreme point of one facet, and therefore, in order to extract Stackelberg from the learner, the optimizer must be able to (or potentially implicitly) reconstruct the facet boundaries around the equilibrium point. We illustrate the definition of facets by the following example:
Example 2.
Let $m=n=3$ and consider the learner payoff matrix $B=I$, with the optimizer action $x=(x_{1},x_{2},x_{3})^{T}$, the facets $E_{1},E_{2}$ and $E_{3}$ are plotted with different colors in $\Delta_{3}$ in Figure 1.
Figure 1: Plot of facets $E_{1},E_{2},E_{3}$.
The dashed lines indicate the boundary between facets.
While similar definitions are made in (Letchford et al., 2009; Peng et al., 2019; Lattimore and Szepesvári, 2019), we use the word facet because each $E_{i}$ induced by any matrix $B$ is indeed a polytope that is a subset of $\Delta_{m}$ and their union $\cup_{i\in[n]}E_{i}=\Delta_{m}$.
@@
At a Stackelberg equilibrium the learner is usually indifferent between multiple pure strategies, which means that switching from the equilibrium pure strategy $y^{*}$ to another indifferent pure strategy $y^{\prime}$ does not incur additional regret to the learner, while potentially vastly degrading the optimizer’s payoff. Therefore, instead of simply selecting the equilibrium point, the optimizer must choose a pessimistic equilibrium point that sacrifices some utility from Stackelberg value to guarantee the learner responds with $y^{*}$. We define pessimistic facets as:
@@
The proof is deferred to Section B.1. As shown in Proposition 1, as long as the optimizer knows a complete set of $d$-pessimistic facets for each $i\in[n]$, it is able to compute an approximate Stackelberg equilibrium up to an error at the scale of $d$. We now extend this result to its ability to extract Stackelberg value against no-regret learners:
@@
A refined version of Theorem 3 that expands the big $O(\cdot)$ notation as well as its proof can be found in Section B.2. Here we can see if we take $d_{1}=d_{2}=\sqrt{f(T)/T}$, we obtain an optimizer Stackelberg regret of $O\left(\sqrt{Tf(T)}\right)$, which is $o(T)$.
@@
To describe an equivalence class $\mathcal{B}$, observe that for all matrices $B$ in $\mathcal{B}$, the matrix $\mathcal{B}_{i}^{\circ}\in\mathbb{R}^{m\times(n-1)}$ defined by $(\mathcal{B}_{i}^{\circ})_{:,k}=(B_{:,k}-B_{:,i})/\max_{j_{1},j_{2}}\|B_{:,j_{1}}-B_{:,j_{2}}\|_{\infty}$ for some $k\neq i$ in each column will be the same for any fixed index $i\in[m]$ with the convention that $/0=0$ 0, so we can use $\mathcal{B}_{i}^{\circ}$ to represent the entire equivalence class. Based on this, we can also define the difference between two equivalence classes:
@@ 5.2 Steering with Payoff Class Estimation
We provide the following example:
Example 4.
Following Example 2, consider
$\hat{B}=\begin{bmatrix}1.05&0.05&0\\ -0.05&1.05&0\\ 0.05&0&0.95\end{bmatrix},$ (22)
for facet $E_{1}$ and the corresponding $\mathcal{B},\hat{\mathcal{B}}$, we have that $\|d_{1}(\mathcal{B},\hat{\mathcal{B}})\|_{\max}=0.1$. we plot $E_{1}^{-}(\hat{B},d)$ and the boundaries of $E_{1}(B),E_{1}(\hat{B})$ in Figure 2.
Figure 2: Comparison among $E_{1}(B),E_{1}(\hat{B})$ and $E_{1}^{-}(\hat{B},d)$.
We can see that although $E_{1}(\hat{B})\nsubseteq E_{1}(B)$, we indeed have $E_{1}^{-}(\hat{B},d)\subseteq E_{1}(B)$.
@@
We can see from Definition 11 that $C_{i}\leq 0$ if and only if (15) is feasible. We show in Section C.2 that if $C_{i}>0$, (15) is infeasible and there exists a gap where it can be relaxed while still being infeasible. If $C_{i}=0$, however, Definition 11 indicates that $\forall x,\max_{j}x^{T}\mathcal{B}_{i}^{\circ}(e_{j}-e_{i})\geq 0$, and $\exists x\in\Delta_{m},j\in[n]$ such that $x^{T}B_{i}(e_{j}-e_{i})=0$. Under this case, the facet $E_{i}(\mathcal{B})$ has zero volume and as long as the estimation $\hat{\mathcal{B}}$ is not precise, the facet $E_{i}^{-}(\hat{\mathcal{B}},d)$ could always be empty. Also, even if the optimizer knows the real underlying $\mathcal{B}$, since $e_{i}$ is weakly dominated, the learner is not steerable if $e_{i}$ happens to be the Stackelberg equilibrium since the learner is indifferent between $e_{i}$ and $e_{j}$. To avoid this special case (which occurs with probability 0 for uniformly randomly generated $B$ matrices (Von Stengel and Zamir, 2010)), we make the following assumption, as is standard (see e.g., (Gan et al., 2023; Deng et al., 2019; Brown et al., 2023)):
@@
Similarly, the refined version of Theorem 6 with explicit Stackelberg regret bound and its proof can be found in Section C.5.
@@ 6 Learning to Steer Classes of Learners
We provide numerical experiments to illustrate the effectiveness of the algorithms in Appendix F.
@@ 6.1 Learning to Steer Ascending Learners with $n=2$
Suppose $m=n=2$ and the payoff matrix $B$ does not contain identical columns. For some chosen parameter $d$, if either one facet is empty, or each facet has diameter at least $d$ and $P_{2}$ uses an ascent algorithm $\mathcal{A}_{2}$ that is $f$-no-regret, Algorithm 1 with accuracy margin $d$ achieves a Stackelberg regret of at most $O(\frac{f(T)}{d}+dT-\log d)$ as long as $d=\Omega(\exp(-f(T)))$.
A more detailed version of Theorem 7 with its proof can be found in Section E.1. As an example, here if $f(T)=T^{\alpha}$ and we take $d=\sqrt{f(T)/T}$, we achieve a bound on the optimizer Stackelberg regret of $O\left(\sqrt{Tf(T)}-\log\sqrt{f(T)/T}\right)=O\left(T^{\frac{1+\alpha}{2}}+\frac{1-\alpha}{2}\log T\right)=O\left(T^{\frac{1+\alpha}{2}}\right)$.